### v1

# Dynamic Neural Curiosity Enhances Learning Flexibility for Autonomous Goal Discovery

Abstract
The autonomous learning of new goals in robotics remains a complex issue to address. Here, we propose a model where curiosity influence learning flexibility. To do so, this paper proposes to root curiosity and attention together by taking inspiration from the Locus Coeruleus-Norepinephrine system along with various cognitive processes such as cognitive persistence and visual habituation. We apply our approach by experimenting with a simulated robotic arm on a set of objects with varying difficulty. The robot first discovers new goals via bottom-up attention through motor babbling with an inhibition of return mechanism, then engage to the learning of goals due to neural activity arising within the curiosity mechanism. The architecture is modelled with dynamic neural fields and the learning of goals such as pushing the objects in diverse directions is supported by the use of forward and inverse models implemented by multi-layer perceptrons. The adoption of dynamic neural fields to model curiosity, habituation and persistence allows the robot to demonstrate various learning trajectories depending on the object. In addition, the approach exhibits interesting properties regarding the learning of similar goals as well as the continuous switch between exploration and exploitation.
Index Terms:
Intrinsic motivation, curiosity, attention, task engagement, locus coeruleus, autonomous goal learning, dynamic neural fields.
I Introduction
Developmental robotics takes inspiration from various fields such as developmental psychology, neuroscience, machine learning or philosophy. A popular domain of interest in the field and one of the core idea of this study is called Curiosity. Curiosity can be seen as a particular case of Intrinsic Motivation (IM) and researchers have demonstrated that IM could be an efficient drive for the autonomous learning of behavior [1]. In addition, the learning of various behavior exhibits a developmental pattern [2, 3]. Nowadays, there are multiple ways of modelling curiosity and one of them consists of training predictors such as forward and inverse models and thus displaying a learning progress that depends on the error variation.
Along with curiosity, attention is a fundamental cognitive process. Indeed, attention allows humans to focus and concentrate on specific elements of our surroundings. However, attention is a broad term gathering several specific mechanisms. For example, visual attention is the ability to sustain the gaze at salient stimuli and is developing during the first months after birth [4], [5]. To help search for specific visual stimuli, inhibition of return is a well studied mechanism naturally preventing a person to look twice at the same location within a short interval of time [6]. In addition, we can distinguish two types of visual attention : exogenous, driven by external stimuli or endogenous and goal-driven [7], [8]. These two types of attention are often referred as bottom-up for the former and top-down for the later, and inhibition of return can happen during both, but whether there are shared neural pathways between them in order to occur is still subject to debate [9]. In the brain, arousal is mediated by the Locus Coeruleus-Norepinephrine (LC-NE) and demonstrates different patterns of neural activity. According to the adaptive gain theory [10], LC neurons exhibit two modes of activation : tonic and phasic. During tonic activation, the subject disengages from the current task and engages in an exploratory behavior. On the contrary, phasic LC activation engages in exploitation and thus is driven by task specific outcomes. It has been demonstrated that IM is involved in tonic and phasic activation of dopamine [11], overlapping certain neural pathways with noradrenaline (norepinephrine) [12] and thus might be directly interacting with the LC-NE system. If curiosity and more especially its learning progress component is related to attention, then the LC-NE system is a potential starting point.
Here, we propose a robotic cognitive architecture for the online discovery and learning of goals where a robotic arm is learning how to interact with simple objects. To do so, we introduce a model situating the LC-NE system at the center of different processes, thus determining if the robot should explore and discover potential goals through bottom-up attention, or exploit and learn to achieve these goals. During both exploration and exploitation, an inhibition of return mechanism generates a set of actions toward the object location until producing a change in the environment. If the inhibition of return is shared by exploration and exploitation, there is however a conceptual difference between them since the discovery of a new goal in the environment rests on exogenous attention, and on a top-down and goal driven approach for the learning of these goals. For the exploration stage, we complete the attentional mechanism connected to the LC-NE system with an habituation paradigm [13], where the discovery of a new goal produces a tonic activation and limits the time of exploration. To balance this process, the absence of a new goal signals a shift from tonic to phasic activation. During the latter, the learning of a goal occurs with the use of simple neural networks implementing a forward and inverse model. The output from these neural networks determine an error associated with a goal and its learning progress. With the exception of the forward and inverse models, the entire architecture is designed with dynamic neural fields [14]. Goals are then represented within neural fields by the neural activation of their learning progress and are projected to the LC-NE system, which then chooses to exploit the goal with the highest learning progress. By design, the model follows an enactive approach by continuously pushing the robot to interact with its environment and thus shaping its cognition by using several forms of working memory. We will demonstrate that when close goals are discovered, the robot autonomously only focuses on the one with the highest error and slowly forgets about the other. We evaluate several neural parameters such as persistence, inhibition of error and habituation that modulate and optimize the learning of goals, depending on the objects. To our knowledge, this is the first attempt to model intrinsic motivation with dynamic neural fields and thus support the robot to express different learning behaviors by tuning intrinsic properties of the dynamic neural fields.
This paper will provide a theoretical background according to the literature in Section II before presenting the architecture of the experiments (Section III). Then, we will present the experimental setup with the frameworks used in Section IV followed by the different results (Section V). Finally, we will discuss about the advantage and limitations of the architecture (Section VI) before concluding.
II Related work
II-A Curiosity and Intrinsic motivation
In terms of biological perspective, intrinsic motivation (IM) needs to be distinguished from extrinsic motivation [15]. The former can be seen as a way for the brain to monitor the learning efficiency of a particular knowledge while the latter directly guides the learning itself through external cues/rewards. Then, there are different metrics to compute intrinsic motivation [16], separating knowledge-based from competence-based systems [17]. There is a consensus in the literature classifying curiosity as a particular case of intrinsic motivation [18]. Indeed, curiosity focuses on the learning progress hypothesis [19], [1] and formulates it as the evolution of prediction error while learning a task. In practice, curiosity can be used for the learning of a world model by leveraging and focusing the sensorimotor experience [20]. GRAIL [21] is a cognitive architecture operating on different levels where the IM used is competence-based, determining the best goal to pursue depending on how well the learning occurs. Intrinsic motivation in robotics is also used to form hypotheses in developmental psychology. For example, a developmental robotics model identified IM to explain the sudden tool-use improvement in children [22]. However, the discovery of goals in visually rich environments is problematic due to the exploration space becoming quickly too vast for an intrinsically motivated system to be efficient. To address this issue, researchers are applying a set of dimension reduction methods. One of the first study to overcome these limitations is the case of a robot arm interacting with a ball [23]. In this case, a variational autoencoder is trained and goals are sampled from the resulting goal space. Other work in the domain and related to robot grasping extended this concept by linking and compressing the object features, the action and the sensori outcomes together [24]. By doing so, the robot creates a latent space where the learning of forward and inverse models is facilitated to achieve grasping. Treating with high dimensional spaces with intrinsic motivation is far from trivial, and the sampling problem can gain insights with a more developmental point of view. In these studies, it is necessary to perform several round of exploration in order to build up the different goal learning progress and then select them for learning. However, it is not clear how the switch between discovery and learning takes place in infancy. In our work, we propose that the error contributes at some extend to ease this shift. Finally, IM proved to be a drive reproducing developmental patterns [25], but actively and efficiently sampling the environment remain difficult to tackle [26]. However, researchers demonstrated that IM involves other cognitive processes, such as attention, and provide perspective toward possible neural basis of intrinsic motivation [27].
II-B Attention and inhibition of return
Visual attention is essential for an individual to focus on an unexpected event and to optimize cognitive processes on a specific task [8]. This distinction separates exogenous (bottom-up) from endogenous (top-down) attention. If motivation can sharpen exogenous spatial attention under certain conditions [28], the modulation of top-down attention can enhance the learning of a particular task [29, 30, 31]. The neural basis for these distinct processes are, for a significant part, shared [32], but not entirely [33]. Regarding the coupling between attention and intrinsic motivation, a study used reinforcement learning to allocate and shift attention [34]. More close to our approach, researchers used bottom-up attention to learn object affordances and decompose a task into sub-goals [35]. Here, we choose to focus on the inhibition of return effect (IOR) to facilitate attention toward an object and ease the discovery of new goals. In the literature, the general mechanism of exploration is often resumed as random and direct exploration [36, 37] where the former consists of producing random actions with a high uncertainty but potentially large expected reward. The latter directs exploration toward more certain rewards but necessitates prior information. With curiosity in robotics, this difference can be translated to a motor babbling behavior for random exploration and the building of representations for direct exploration. For example, a robot can learn a representation of a diverse set of goals encountered during random exploration to generate new goals [24]. Here, our purpose is to propose an additional exploration method that can reduce the uncertainty of random exploration via the inhibition of return effect. This mechanism was pointed out by Posner [6] in visual attention when he discovered that the return of attention to a cue previously attended expresses a longer time. On the opposite, there is a facilitation (faster response time) if the time interval between cue-targets is short. This consequence endows the brain with a natural foraging mechanism of visual stimuli [38], but it remains necessary to proceed to an attentional disengagement in order to observe the IOR effect [39]. However, it appears that an endogenous disengagement of attention is not enough in order to observe the effect and is instead always relying on exogenous disengagement of attention [40]. In addition, it is possible to induce an inhibitory of return for both exogenous (uncued location) [41], [9] and endogenous (cued location) attention, but the effect seems highly correlated to the cue’s saliency and thus relies on an important exogenous component [42]. Regarding IOR in action production, there is no consensus whether the effect is directly or partially involved. Several research reproduced reaction time of IOR during a hand reaching movement [43] and even showed that reaching paths were biased toward the cue locations under certain conditions [44]. Directly connected to our research, it has been determined that IOR effects can be applied to objects [45]. Even more, inhibition can interact with working memory to mediate goal-directed action toward an object, optimizing performance for future interaction. In consequence, we intend to generate and use IOR around the object location so the robot can generate goal-driven actions. To be able to identify these two forms of attention and decide to engage in exploration, we take inspiration from the Locus Coeruleus.
II-C Exploration, exploitation and the Locus Coeruleus system
The locus coeruleus-norepinephrine system (LC-NE) is a brainstem nuclei that is widely connected to the neocortex. The first unifying approach regarding its role and functions is the adaptive gain theory [10]. The theory demonstrates two different neural activations, depending on the arousal state as well as the relation with task engagement. More precisely, the LC-NE system exhibits a tonic neural activation during exploration, rising arousal’s level and thus bottom-up attention toward salient events [46]. Here, the exploration strategy consists of generating goal poses around the object with IOR and observe the possible outcomes. A phasic activation suggests a goal-oriented behavior (exploitation) with a focus on the task (top-down attention). Then, the LC-NE system is a crossroad between attentional processes and decision making. Aston-Jones and Cohen determine the emergence of a tonic activation when the utility of a task vanishes. The LC-NE receives direct connection from the anterior cingulate (Acc) and orbitofrontal cortices (OFC) which are directly involved in task-related utility and decision making [47], [48]. More specifically, the ACC seems to represent value predictions about reward and, more generally, uncertainty [49]. There is a large consensus to recognize the basal ganglia as the processing place of rewards that can support reinforcement learning [50]. Several robotics experiments take inspiration from the phasic delivery of dopamine to signal rewards based on prediction errors [51], [52]. We can observe a similarity between the phasic delivery of noradrenaline from the LC and the phasic dopamine-based signal emerging from the basal ganglia, especially after pointing out that dopamine and noradrenaline share neural pathways [53]. In addition, researchers demonstrate that if both dopamine and noradrenaline are related to rewards prediction, only the latter is predictive of task engagement and is strongly activated when cues indicate a new task condition. This means that on the contrary to dopamine, noradrenaline is strongly involved in motivating current or future engagement. The Acc directly projects prediction information about the task utility to the LC-NE system which in turn decides if the utility is high enough to drive the neural activation in a phasic mode, hence to an exploitatory stage. Following the predictive coding approach from Friston [54], researchers proposed a model of active inference base on the LC-NE system’s ability to track prediction errors by linking states and actions together [55]. The interesting point of the study is how the LC can possibly rely on a prediction error mechanism and thus be related to intrinsic motivation. Under the prism of active inference, they demonstrate that the two firing modes of the LC are an emergent property from the processing of state-action prediction errors. Furthermore, scientists used functional magnetic resonance imaging (fMRI) to determine the neural substrates of intrinsic motivation [56]. They suggest that diminished patterns of neural activity correlate with individuals exhibiting high intrinsic motivation. This finding does not specify which brain region is concerned by intrinsic motivation, yet high intrinsic motivation seems to rely on a phasic neural activation. Therefore, we believe that the functioning of the LC system is promising when using curiosity for robotics. It will allow the robot to engage in exploration and exploitation depending on uncertainty, meaning on how accurate the robot’s predictions are, based on the learning progress. As a consequence, the LC model will follow a tonic or phasic neural activation, thus autonomously tuning attention to modulate the level of arousal and engage in exploratory activity or the learning of a skill.
II-D Habituation and Persistence
In order to limit the time spent on exploration, our approach introduces a cognitive process called habituation. Indeed, there is no consensus as to how the LC switch from a tonic mode with high arousal and engages exploration to a phasic mode (exploitation). Here, we propose to model that switch by an habituation paradigm where the phasic activation from LC happens only when the robot’s perception of an object does not produce a novelty effect anymore. Habituation is well studied by researchers, especially in developmental psychology by evaluating how long infants examine a new stimuli. Research first focused on the link between familiarity of the perception and exploratory behavior [57] and concluded that infants that have been familiarized to some objects prefer to look at new ones. A similar study [58] confirms that infants of seven and twelve months old have a decrease of attention when the objects become familiar. Interestingly, the study determined that the looking duration was not dependent on the age but more on the familiarity of the object. If these findings concern an exploration behavior and how habituation can influence it, the duration at which an individual learns a new skill must be taken into account.
Cognitive persistence is a fundamental aptitude if one wants to learn a new goal. Observing how individuals persevere to learn something new can help to decide how and if a learning is actually fruitful. Regarding developmental psychology, the measure of persistence is investigated by assigning tasks to infants and evaluating how long they engage in a learning activity [59]. Here, researchers established that there is a link between an infant’s persistence and competence. This means that infants who were more stimulated at home persist more toward goal oriented actions. On the opposite, infants with cognitive delay demonstrated less persistence, as well as engaging toward less challenging learning activities [60]. From a neuroscience perspective, there is little known about cognitive persistence. Teubner-Rhodes [61] investigated this ability to overcome task difficulty in adults. With the use of fMRI, she located persistence in the prefrontal cortex and concluded that adults with high perseverance demonstrated better performance and adaptive control for the task. Later, a study focused on persistence for multilingual people [62] and located this process in the inferior frontal gyrus and dorsal anterior cingulate cortex (dACC). This could suggest that in addition to dealing with uncertainty and prediction-error, ACC is also involved in cognitive persistence.
These cognitive processes are necessary for our approach in order to support the shift from tonic to phasic activation for habituation, and to determine how persistent the robot has to be while learning a skill. Each aspect of cognition described in this section will be modelled with the Dynamic Field Theory.
II-E Dynamic Field Theory
In this work, we model the different cognitive process with neural dynamics and more especially Dynamic Field Theory (DFT) [14]. By doing so, the system exhibits a range of different behaviors by only tuning the intrinsic parameters of the neural fields. The theoretical framework mathematically models the evolution and activity in time of large populations of neurons. The approach has been successful to model complex cognitive processed, such as visual working memory [63], intentionality [64] or motor habituation [65]. The formulation and implementation of DFT is provided in the Appendix -A. Regarding the modelling of motions, we adopted the dynamic movement primitive framework.
II-F Dynamic Movement Primitive
In this work, we control the motion of a robotic arm by using the Dynamic Movement Primitive framework (DMP) [66], as they can exhibit a wide range of motions with a limited number of parameters [67]. The approach uses a set of differential equations to represent a movement with the advantage of being robust against perturbation. More precisely, we adopt the formalism introduced by Pastor [68], which allows to reach the generalization of a motion by adapting a start and goal parameter. More details about the method and its implementation can be found in the appendix -B.
III Architecture
III-A Overview
Figure 1: (Left) Architecture flow from goal discovery with bottom-up attention (new event) to learning with curiosity. (Right) Connectivity flow between all components in terms of exploration (goal discovery) and exploitation (goal learning).
The Locus Coeruleus (LC) architecture is the main component of our model. In the brain, this nucleus is directly responsible for attention and task commitment. Here, we link intrinsic motivation and the LC activity through the engagement in learning new goals with different types of attention. We propose to articulate and describe how these cognitive processes might work together in this section.
The LC exhibits two modes of neural activation : tonic for exogenous attention and a phasic spiking mode for top-down attention during task engagement. Because of the time scale abstraction level, the DFT framework can not precisely differentiate them as well as biological neurons. Then, we define a tonic activation as a sustained and uninterrupted excitation activity while the phasic mode exhibits a more sparse and localized activation in time. Figure 1-left depicts the architecture flow from goal discovery with bottom-up attention to learning with curiosity, followed by the robot. At first, the robot detects any salient stimuli that can take place in the camera field of view. Then, the process of goal discovery begins by performing action babbling around the object. During this stage, we can consider that the type of attention is exogenous, since the robot is sensitive to any stimulis (objects) in the environment. At the same time, an habituation mechanism reduces the sensitivity towards the objects, thus reducing the neural activity within the LC. Then, the robot selects a goal to learn with the help of the curiosity and persistence mechanisms. At that step, there is a shift from bottom-up to top-down attention since the model only focuses on observing outcomes from a particular object. Finally, the robot can switch attention to a new object and starts over the process of goal discovery. However, this last step could be ignored and the robot will choose to discover new goals on a previously seen object.
In this work, we define exploration as the process of goal discovery and exploitation as the stage to learn these goals (See Fig. 1-right). The perception and action module are involved in both steps by discovering new goals and processing already seen stimulis for the former, and generating or using a DMP (dynamic motion primitive) for the latter [66]. The perception system observes the outcome of an action by coupling the color of the object along with its motion in space. Then, the habituation module is in charge to determine if the observed stimulis is seen as new or if the robot is already habituated to this specific object. The persistence, error and learning progress module are taking place during exploitation. Persistence determines for how long the robot should attempt the learning of the current goal. The error module monitors the discrepancy between the output from the forward model and the actual perceptual outcome, hence supporting the neural computation of the learning progress. During both stages, the perception engine sends direct inputs to the forward and inverse models present in the error component. The following section presents details the Locus-Coeruleus model.
III-B Locus Coeruleus
Figure 2: Overview of the Locus Coeruleus component driving goal discovery and learning.
The LC mechanism is coordinating the engagement level regarding the discovery or the learning of goals (See Fig. 2). At any time, the active node always exhibits a supra threshold level of activation and excites the slow boost component, a particular form of memory trace that evolves depending on two nodes inputs (Appendix -A). This memory trace builds activation when the active node projects excitation, maintains the current level if both connected nodes are down, and decays the activation if only the threshold node is up. At the beginning, the slow boost module slowly rises the resting level of the phasic and tonic neural field (NF) and stops when a peak of activation appears, triggering the stop node. The tonic activity of neurons is depicted with the tonic NF and the excitation incoming from the habituation mechanism (scene and object selection NF). This neural field exhibits a supra-threshold peak of activation as long as the robot is not accustomed to an object. The phasic NF gathers activations from the error and learning progress models but with a higher resting level. This means that after discovering goals, the phasic NF will be the first to see the emergence of supra-threshold activations. Tonic NF has a resting level of -2 and phasic NF -1.05. This implies that the LC can see a peak in tonic NF when the cumulated learning progress and error are below 0.05. This indicates that the LC can drive into exploration again when all the skills are learned. After discovering several goals through bottom-up attention and being habituated to an object, the robot does not exhibit any learning progress. Then, it is necessary to bind the errors memory trace to the phasic NF in order to bootstrap the learning of previously discovered goals. We empirically chose the gain factor so the robot can engage in learning a skill if there is no learning progress ($\approx$ 30 percents of the errors present in the error module). However, we keep the errors contribution low enough to favor the learning of a task that has a small learning progress. By doing so, the robot can begin the learning of goals that do not exhibit any learning progress instead of performing several stage of exploitation for each goal beforehand to build up the different learning progresses. The one dimensional object decision NF represents the object’s color in which the robot should explore or engage in learning.
By analyzing the connectivity flow of the model, we can see that we prioritize exploration the same manner that exogenous attention seems to be operating within the LC. As mentioned in the literature (II), the engagement in a task can only operate if there are no salient stimuli and thus a low level of bottom-up attention. The priority to exploration comes along with the tonic activation. Our approach demonstrates a sustained activation (i.e tonic) while discovering new goals, thus preventing any possibility to learn a specific skill. During exploitation, the model selects a goal to learn on a regular basis. This process stops any activation within the phasic NF and raises again the resting level to select the goal with the highest learning progress (or error in the absence of any learning progress). In this model, the persistence mechanism directly influence the goal selection and depict a phasic activity. Finally, if the robot is habituated and has learned all the goals discovered, the tonic NF will select a know object to perform exploration again (Appendix -C).
III-C Perception and habituation
The perception module processes the object’s motion direction after an action. A perception is formalized by a 2D neural field, where the horizontal and vertical dimensions represent the object color and motion’s angle as a goal, respectively (see Appendix -E for more details).
Figure 3: Architecture of the perception mechanism. It gathers the goal detector and the Hebbian learning between a goal and a DMP.
The motion of the object is defined by an angle between a static reference vector and the observed vector’s motion of the object. The angle is bound between -$\pi$ and $\pi$ radians and scaled between [0,100] to fit the neural field feature space. It is important to emphasize that the detection of the object’s motion takes place in real-time, i.e., while the robot performs the action), delivering a sustained input to the goal perceived NF (Neural Field). In Fig. 3, an inhibition of return is taking place with goal perceived NF, WM goals and the goals NF. This avoid perceiving a stimuli as new when the goal has already been discovered. During exploration and if the stimuli is new, a dynamic motion primitive (DMP) is generated by the robot controller and activates the associated DMP neuron through a one-to-one connection. The Hebbian learning scheme associates the generated DMP with the perceived outcome (goal focus NF) with a reward-gated rule similar to [64] :
$\begin{split}\dot{w}_{{dmp}_{i},{col,ang}}=-\eta\,\sigma(LC:explore)\,\sigma(dmp_{i})\\ ({w}_{{dmp}_{i},{col,ang}}-\sigma(g(col,ang)))\end{split}$ (1)
The learning occurs when there are activations from DMP neuron, learning signal and the goal focus NF. They are respectively denoted by $\sigma(dmp_{i})$, $\sigma(lc:explore)$ and $\sigma(g(col,ang))$. The term $\sigma(g(col,ang))$ represents the sigmoid activation of the goal focus NF at the object color ($col$) and angle ($ang$) location. When the robot engages in learning a task (i.e., exploitation) the goals NF is inhibited so a peak of activation can emerge within the goal focus neural field. Then, the activation spreads through the synaptic weights and triggers the DMP neuron corresponding to a particular goal.
Figure 4: Neural flow of the habituation process decreasing the objects novelty in time.
The habituation mechanism is tightly linked with the perception module and is inspired by the work of Perone and Spencer [69]. The mechanism demonstrates how infants become familiar to similar stimuli in time and lead to a decrease of novelty. Here, the stimuli are the color and the motion features (i.e., within the perception module). The former processes input given by the latter and drives the LC from a tonic to a phasic mode. The color of the object serves as feature within the object selection NF (Fig. 4). The new perception field receives activation from the goal color dimension as well as from the explore node inside the LC module, supporting a supra-threshold activation only during exploration. The condition of satisfaction node (CoS EOA) is only active for a brief period of time when the robot has finished an action. This node correspond to the term a(t) in Eqn. (4). When a new goal occurs in the scene, the visual memory slowly decays at the color feature location, diminishing the inhibition to the object selection field. On the contrary, if an action results without the perception of a new goal, the visual memory builds up an activation that will inhibit the object selection. In practice, this implies that the robot keeps the exploration of new goals as long as the inhibition is weak enough within the object selection NF. The main difference with the habituation model of Perone and Spencer rest in the coupling between WM colors and the visual memory trace. Here, the inhibition of the visual memory trace is directly linked to the WM colors while their model associate an inhibitory layer with an Hebbian mechanism to the object selection NF. In our model, it results with the object remaining habituated even if the object is out of sight (supra threshold activation within WM colors) and with the possibility to decay inhibition when a new stimuli is observed.
Here, the use of dynamic neural fields provides the possibility to design a more robust perception system. The inhibition of return prevents to identify a previously discovered goal as a new one if it is seen again during exploration. By tuning the neural field parameters (i.e the standard deviation of peaks), it is possible to adjust the likelihood of two goals. With a high standard deviation, two similar goals have a lower probability to be distinguished from each other. In addition, dynamic field theory demonstrated excellent results in modelling brain processes as habituation. This is to our knowledge the first curiosity driven architecture that integrates such mechanism to favor exploration toward novel objects. The following section will describe the action and the persistence system.
III-D Action Formation
Figure 5: Overview of the action module. The model employs an inhibition of return mechanism to generate two poses during goal discovery. Only a single pose is produced and used with a DMP while learning.
The role of this process is to perform action babbling around the object located in the scene (Fig. 5). During exploration, the module generates two poses ($x_{1}$,$y_{1}$) and ($x_{goal}$,$y_{goal}$) that the end-effector will go through from a resting position ($x_{start}$,$y_{start}$). In this work, the z dimension is fixed at one centimeter above the table for a generated pose. The robot records the end-effector position from the resting pose until reaching the second pose and creates a DMP if it leads to the discovery of a new stimuli. While learning a skill (exploitation), the robot activates the DMP parameters of a goal and only needs to create a single pose.
Related to the neural fields flow, a peak of activation emerges within the position object NF when an object is in the camera field of view. This neural field along with the other two dimensional fields present in the module are representing the x and y position of the object in the robot space. Then, we apply a specific convolution between the position object output and the action formation NF. The result of this convolution serves as inhibition to create poses necessary for the DMP creation/activation. The input to the action formation NF can be seen as a reversed Mexican-hat shape centered at the object location (see Fig. 6.
Figure 6: (Left) Mexican-hat convolution with inhibition applied to the object (input to action formation NF). (Middle) Generated pose after rising the resting level within the action formation NF. (Right) The input peak position provided to the robot controller (filtering NF).
In order to generate a pose, we use a slow boost component as seen in the LC module (see Section III-B). The slow boost builds up an activation if signal action is active, and the activation decays when the robot signals the end of an action (robot controller). The reversed Mexican-hat shape input to action formation favors peaks of activation around the object location. The signal action node receives inhibition when the robot is moving and when a peak emerges within action formation. The working memory keeps track of the activations in time and projects the inhibition back to action formation. Together, these two fields provide an inhibition of return effect, avoiding to select the same pose twice. Here, we provide a specific value for the global inhibition of the working memory (Appendix -A), so several supra-threshold activations can take place at the same time without over population. This means that only a limited number of activations can be self-sustained. If a new peak emerges, one of the ”old” activations dissipates. The robot controller is in charge of delivering excitation after two poses during exploration and after a single pose for exploitation. The working memory receives inhibition in case of a new stimuli (transient goal) and when the persistence mechanism reached the threshold.
Figure 7: General view of the persistence component. The module adapts the time spend on learning a goal depending on the error level of the forward model
The persistence module bears a rhythm to the action module by signaling when to reset the inhibition of return while learning a specific goal. In the case where the goal error does not decrease, this module stops any further trial to learn a particular goal and clear the memory from all the generated pose. The core of the process lays with the threshold, force and action nodes. Threshold node is a classic zero dimensional neural field, transient action and force nodes are memory traces with a slow build up and fast decay for the former, and a quick build up and long decay for the latter. Threshold node receives inhibition from the local error level (i.e the error of the current goal) and excitation from the transient action. However, the current goal NF output is either zero or one because of the use of the absolute sigmoid as kernel activation. The output from the current goal is always higher than the local goal error who is bounded by the hyperbolic tangent function (section III-E). After each motion, the activation level within transient action gradually increases and spreads to the threshold node. Once the level of transient action overcomes the inhibition of the local error, it excites the attempts memory trace (a(t) term in Eqn. (4)) and leads to an update of the attempts memory trace. To resume, no activation happens within the attempts MT at the beginning of learning, meaning that there is no inhibition toward the LC:phasic. After several trials and if the error remains high, the inhibition will prevent the LC from pursuing the learning of that particular goal in the immediate future. However, the inhibition of a particular goal decreases after focusing on a different skill. By tuning the build up time ($\tau_{+}$) of transient action, it is possible to modulate the number of motion attempts. If $\tau_{+}$ is large, the robot will significantly increase the number of trial to learn a goal. By setting a high value for $\tau_{-}$ (attempts MT), the error level will decay slower. This would extend the time scale in which a goal can be selected again for a learning phase and a high goal error can expand it even further. Finally, the force node is triggered under the excitation of the threshold node and will inhibits the IOR and the activation node in LC to avoid rising the resting level of the slow boost component within the LC. This last inhibition essentially delay for a short time the next goal selection while the IOR activations within the action module are not completely down.
Regarding the action module, one of the technical novelty rests in the use of the convolution applied to the object as well as the inhibition of return. This agency provides a simple and efficient way to generate poses around the object and provide a mean to focus the random exploration [36]. As for the persistence mechanism, a key novelty is to provide a mechanism that could adapt depending on the learning behavior. The use of dynamic neural fields support the flexibility to modulate the number of attempts made by the robot as well as the inhibition of the goal for future selection if the error remains significant. As a result, it is possible to observe diverse learning behavior by tuning the parameters of these neural fields (see sec. V-C)
III-E Curiosity
Error module
When the robot discovers a new goal, a forward and an inverse model are created. The forward model output the first error in order to compute the learning progress with dynamic neural fields (Fig. 8). During exploitation, the robot performs a motion and observes the outcomes in the environment. If it results to no changes (the object did not move), then the error remains the same. In case of any changes, the error controller computes the error with the forward model’s prediction. Here, the model follows a precise timing for the error computation where the observation takes place ahead of the end of the motion. In practice, this mean the robot pushes the object, computes the error while going back to a resting position and finally signals that the motion is done. The curiosity mechanism is then divided into two parts : the error and learning progress processes. The former focuses on characterizing the error of a goal as neural activity while the latter computes the error’s level dynamic to deliver a learning progress. We will first focus on the error architecture before presenting the learning progress structure.
Figure 8: Architecture of the error component. The process formalizes the error coming from the forward model of a specific goal as neural activation.
The compute the error NF gathers error levels from all the goals discovered. When the robot pursues the learning of a skill, the error level from goal error MT is combined with the filter MT in order to only affect the current goal. Indeed, this last field regroups each error and inhibits the current goal whereas the goal error MT provides a single peak (i.e., the goal error’s activation). It is worth to mention that the activation function ($\sigma$) used for this field, local goal NF, local error NF and updated error NF is a ReLU function. The time node acts as a trigger when a new error value is computed. Once a new error is delivered by the error controller, this brief activation supports the update of the memory trace as well as boosting the resting level of other neural fields. The errors MT is there to keep in memory each goal value and only updates them once the end of the action node (EOA) is active. The last main fields are the local goal NF and updated error NF. They have a resting level of $-2$, meaning that they yield peaks of activation only when they receive the boost from the time node. As mentioned earlier, the error is computed before the robot reaches its resting pose and activates the EOA node. This implies that when the time node is active, the local error NF contains the error’s goal before the update whereas the updated error NF delivers the newly calculated value. These last fields prepare the future neural computation of the various learning progresses.
Figure 9: Neural flow of the learning progress. The module processes neural dynamics coming from the error component to generate a learning progress.
Learning progress component
The learning progress mechanism is straightforward (Fig. 9). It receives inputs from updated error NF and local goal NF that are projected into zero dimensional nodes with resting level of $-1.025$. The time node is the same seen in the error architecture. With this pattern of incoming activation, the rise and low nodes are responsible for detecting if the updated error value is respectively superior or inferior to the previous level of activation (local goal NF). The filter NF follows the same purpose as the one described in the error architecture, providing the error level of each goal except the current one. Then, the compute LP memory trace is where the learning progress of a goal is computed. In the event where the low node is active, the supra-threshold activation from variation NF will project into compute LP and build up an activation. When the rise node is active, it indicates that the error level is rising. In that case, there will be no activation within variation NF and compute LP will decay the learning progress of the current goal. If an activation was present within variation NF, the activation will rise depending on the value of ($\tau_{+}$). Finally, there is a case where the values of the error before and after update are too close to each other, which signify that the progress is not significant. Then, the rise and low nodes will not activate. The specific resting level value of these nodes support the fine tuning of this specific event. Here, there will be no activation if the error does not vary beyond a threshold of 0.025.
IV Experimental setup
The purpose of the experiments is to evaluate the learning behavior of the architecture under different parameters and with different objects. More specifically, we want to determine the role of habituation, persistence and error inhibition during goal discovery and learning. We will also analyse the tonic and phasic activation within the LC. Then, we will assess how the Locus Coeruleus model gathering bottom-up attention and curiosity lead to an open-ended learning approach.
Figure 10: (Left) The overall scene of the experiment setup with the Franka Panda in a resting position. The polygon hovering over the table is the virtual camera. (Right) Field of view of the camera with the red ball in sight.
The scene simulates a Franka Emika collaborative robot in Gazebo (Fig. 10). The dynamic neural fields are simulated with the software Cedar [70] and the middleware ROS [71] insures the communication. The objects used for the setup are a red ball, a blue cube and a yellow cylinder. We specifically choose these objects in regard to their difficulty to be manipulated. Since the end-effector orientation is fixed with the gripper closed, it is relatively easy to be able to push the cube in different directions. For the cylinder, the last pose generated has to be more precise to avoid seeing the object slip away during a motion. The ball is the most difficult object to learn to manipulate since a slight error in the end effector position can push the ball in various direction. The experiment begins when an object is present in the camera visual field. The object is respawned at the center of the table if it is out of reach from the robot. It means that the objects can be moved and rotated from various location on the table as long as it is not out of reach. Each neural field is bounded between 0 and 100 along its respective dimension. The gripper orientation is fixed as well as the z dimension while performing a motion. During the goal discovery stage, the end-effector position of the robot is recorded from the resting state until reaching the goal pose (III). The number of recorded poses remains almost the same, depending on the motion ($\approx$ 30 poses) and they are used to generate a DMP if the motion led to a new stimuli.
Perception
The visual perception is processed by a virtual camera with a resolution of 1280$\times$720. This camera is calibrated to the robot work space in order to transform the position of an object from the camera point of view to the robot base. A tracker identifies the object by color and returns its position to the position object NF in the action module. At the same time, the color tracker returns a continuous input to several neural fields (perception and habituation modules). As mentioned earlier, a goal is formalized within a neural field by collecting the motion direction and color of the object along the horizontal and vertical dimensions respectively.
Learning
The learning of a skill is formalized by a forward and inverse model through multi-layer perceptrons. The forward model has one input layer (4D), one hidden layer (6D) and one output layer (1D). The inputs are the current position of the object (x,y) and the motor command (final pose of the end-effector $x,y$). Then, the output of this neural network is the motion (goal angle) of the object. The inverse model has one layer input (3D) composed of the current object position (x,y) and the goal angle of the object. After the hidden layer (4D), the output (2D) is the motor command (end-effector position on x and y). Each DMP is associated with its own forward and inverse model. The purpose of the inverse model associated with a DMP is to learn to generate the correct end-effector final pose for a given goal and object position. More details of the learning parameters can be found in Appendix -F.
V Results
V-A Habituation
In this section, we evaluate the effect of habituation during exploration (see Fig. 11). To do so, we performed ten explorations for each of the three objects (cube, cylinder and ball) by varying the $\tau_{+}$ (Eq. 4) parameter of the visual memory trace.
Figure 11: (Top) Average number of goals discovered for 10 experiments on each object with fast and slow habituation. The values for fast and slow habituation ($\tau_{+}$) are 2 and 4 seconds.
As expected, a slow habituation ($\tau_{+}=4$ seconds) allows the robot to spend more time on exploration, leading to an increase in the number of goals discovered (see Fig. 11a). We can notice a difference among the objects with the ratio :
$R_{avg}=S_{avg}/F_{avg}$ (2)
Figure 12: Neural activation representing the evolution of the learning progress in time with standard deviation across 20 experiments for the baseline experiment. The figure depicts the average learning progress for the cube (left), the cylinder (middle) and the ball (right).
where $S_{avg}$ and $F_{avg}$ are the average number of goals discovered for a slow and fast habituation. The ratio for the cube, cylinder and the ball are respectively 1.53, 2.04 and 1.69. This suggest that given enough habituation time, the robot can double the number of goals discovered with the cylinder. This can be explained by the orientation of the gripper and the inhibition of return. Since the gripper’s orientation is fixed, only a limited number of actions can trigger the discovery of a new goal. The movement is guided by the inhibition of return around the object’s location, so this inhibition has to be precise and thus requires more time to generate a meaningful motion. In the following section, we will analyse the learning trajectory of different goals with curiosity.
V-B Goal learning progress
Baseline Error Inhibition Persistence
Skill sequence Cube Cylinder Ball Cube Cylinder Ball Cube Cylinder Ball
up $\rightarrow$ left 3.15e-5 0.03 1.94e-6 0.1696 1.27e-3 1.1e-4 2.83e-5 3.84e-6 1.21e-5
left $\rightarrow$ right 3.85e-9 1.85e-7 1.94e-6 2.21e-4 1.27e-3 1.33e-3 2.83e-5 3.84e-6 2.12e-4
right $\rightarrow$ down 1.18e-9 5.73e-7 2.69e-6 2.38e-4 1.99e-3 7.98e-4 3.21e-5 4.3e-6 03.14e-4
Table I: Mann-Whitney U test between 2 consecutive skills for each object, under each setting and for all experiments.
V-B1 Time-course of learning progress between objects
The evolution of the learning progress is presented in Fig. 12. We run the experiment twenty times for each object and, to compare the learning between objects and parameters, we fixed the same four goals. These goals are : push up, push to the left, push to the right, push down. The weights initialization of the forward models determine the first error during goal discovery, which then influence the future goal selection for learning. We consider this experiment as the baseline, with a persistence value (transient action MT node) at $4500ms$ and $100ms$ for respectively $\tau_{+}$ and $\tau_{-}$. The error inhibition (attempts MT projected on LC:phasic) is settled at $100ms$ for $\tau_{+}$ and $\tau_{-}$, which means there is no error inhibition staying in memory between the selection of a goal to pursue. The learning progress (LP MT in learning progress module) is fixed at $2000ms$ for $\tau_{-}$ and $\tau_{+}$. Regarding the LP, we arbitrarily choose these values because a high $\tau_{-}$ would signify that the goal error takes more time to decay even if the error coming from the forward model is already steady at a low level. We performed a Kruskal-Wallis test to evaluate the significance of the skills ordering for each object. For the cube, we took the distribution from $t=[0,1800]s$; For the cylinder, we sample $t=[0,2250]s$; For the ball, the distribution range is $t=[0,6000]s$. The differences of distribution between the four goals are significant with $H(4)$, $p<0.01$ among the three objects. Even if the sequence of the four goals is significant among all the runs, we still perform a Mann-Whitney U test (see Tab. I) between two successive goals (push-up $\rightarrow$ push-left, push-left $\rightarrow$ push-right, push-right $\rightarrow$ push-down) to verify a possible overlap between them in certain runs. The statistical test for the cylinder is almost not significant, which indicates a different ordering between $up$ and $left$ among several runs. We can conclude that the goals follow the same order of learning for the baseline experiment. However, these skills do not reflect a developmental trajectory since they can be learned independently. Moreover, the learning begins with the goal that has the highest error, which depends on the synaptic weights initialization of the forward model. The difficulty of a goal to be learned can be evaluated by the time spent until the learning progress activation comes below $0.01$. For the ball, it is considerably more difficult to learn the set of goals than the cylinder and the cube ($\approx$ 6000s). This is essentially due to the precision required to generate a $goal$ (Appendix -B).
V-B2 Winner inhibits similar goals
Figure 13: (Left) Three goal errors in errors MT belonging to the same object where two of them are similar. The highest peak is selected as current goal for learning when there are no learning progress. (Right) Effect of the inhibition of the current goal on its neighborhood inside filter MT.
During the exploration with the habituation phase, it is possible for the robot to discover similar goals. For example, the robot can push a ball up and push a ball up but slightly shifted by a few degrees. In that case, only the goal with the highest error coming from the forward model will be learned in the future. This comes from the neural dynamics of the filter MT (Error module), and more particularly from the projection of the current goal (see Fig. 13). Indeed, the filter MT represents the various goal errors discovered except the goal being currently learned via the inhibition of current goal NF. This localized inhibition not only affects the goal peak, but also partially or completely alters any peaks of activation at proximity. This effect depends on the standard deviation ($\sigma$) of the inhibition. A nearby goal will be affected by having its peak level lowered, then updated within the compute error NF. This essentially indicates that each time the error is computed for the current goal, any adjacent goals will see a fall on their activation level. A higher standard deviation for the peak activation or a wider inhibition value from current goal NF means that the goals must significantly differ from each other. Indeed, it can be seen as a way to first focus on acquiring a diverse set of simple skills before specializing. The consequences of this effect lead to developmental questions and should be investigated more deeply.
Figure 14: Top row : goals learning progress across 20 experiments under the error inhibition parameters for the cube (left), the cylinder (middle) and the ball (right). Second row : goals learning progress across 20 experiments under the persistence parameters for the cube (left), the cylinder (middle) and the ball (right).
V-C Influence of inhibition and persistence
After the learning of goals under the baseline parameters, we are adjusting the error inhibition dynamics that will be projected to the LC (via attempts MT). We settle the time parameters at $\tau_{+}=2000ms$ and $\tau_{-}=1500ms$ and run the experiment twenty times for each object (see Fig. 14, top row). We arbitrarily chose these parameters so the inhibition takes more time to build up and decay. For a goal where the error is difficult to decrease, it will deliver a mild inhibition that would avoid the goal to be selected again in a near future. We repeated a Kruskal-Wallis test for the cube, cylinder and ball for the time intervals $t=[0,2100]s$, $t=[0,2500]s$ and $t=[0,4250]s$ respectively. The distribution of the four skills across the 20 experiments was conclusive with $H(4),p<0.01$. A Mann-Whitney test was performed to investigate if the goals are following the same order consecutively (see Table I). We observed that the test is not significant for the cube, which signifies that there is no clear ordering between $pushing-up$ and $pushing-left$. Indeed, the $pushing-up$ goal being difficult to learn, if the learning progress did not rise after a run of learning, the inhibition will prevent the selection of this goal in the near future. The statistical test result was significant for the cylinder and the ball but with a higher value than the baseline experiment. This demonstrates the presence of several runs for both objects where the goals do not follow the mainstream sequence pattern. If we compare the inhibition settings with baseline, we can see a rise in the learning time for the cube and cylinder. However, this learning time is shorter for the ball and the goals $push-right$ and $push-down$ are less subject to fluctuation. For the ball baseline experiment, $push-right$ and $push-down$ take respectively $5000s$ and $3900s$ against $3000s$ and $2000s$ under the inhibition parameters. The baseline experiment only relies on the learning progress of the goals to continue the exploitation, which provokes only minimal learning fluctuation in case of easy objects to learn (cube and cylinder). For a difficult object, the system has to continue the learning of hard goals that exhibits a certain learning progress with no possibility to select a different skill to exploit. Under the inhibition parameters, the model introduces more flexibility in the goal exploitation since the learning progress is not the only factor of selection. Essentially, this suggests that the robot does not finish learning a goal entirely before learning a new one. This switch between goals indicates an advantage on the ball whereas it introduces more fluctuation for the cube. Apart from a certain learning time extension ($\approx 250s$), the cylinder does not seem to be affected by the inhibition parameters.
We introduce the persistence settings by modifying the transient action MT node ($\tau_{+}=6000ms$, $\tau_{-}=100ms$). These settings are extending the time of exploitation before the LC module updates its decision regarding the goal to learn. This parameter is analog to the exploration time seen in studies implementing intrinsically motivated systems [22], [24]. The experiment was repeated twenty times for each object and we performed the Kruskal-Wallis test to analyse the goal distributions. The test was conclusive for the three objects with $H(4),p<0.01$ and the Mann-Whitney U test between two consecutive goals depicts no significant variation in the learning sequence for all of them (see Table I). However, the statistical signification is closer to the baseline experiment than for the inhibition parameters. The persistence parameters provide similar results to baseline for the cube. Regarding the cylinder, the distributions of the different learning progress are more distinct. This manifests less learning overlap between goals across the runs. Regarding the ball, the learning time for all skills is reduced compared to baseline but increased compared to inhibition.
Experimenting on three different objects with various learning complexity allows to evaluate the error inhibition and persistence. Regarding the former, this neural mechanism is introduced to support the switch of goal learning if the learning progress does not evolve. For the latter, we demonstrate that the time window spent on trying to learn a goal can have an influence depending on the object. Together, these parameters demonstrate that they could improve the learning of different skills by relying on the complexity of the objects. Here, the baseline experiment is optimal for the cube with no significant learning fluctuation. The ball exhibits the shortest learning time with moderate learning fluctuation under the inhibition settings and the cylinder displays a weaker overlap of learning between goals under the persistence criterion.
V-D Tonic and phasic activation
Figure 15: Time course of the tonic and phasic activation for the discovery and learning of a single goal.
The slow boost component is the core of the tonic and phasic activation within the Locus Coeruleus model. In Figure 15, we monitored the activations within the tonic and phasic NF. During exploration, the tonic activation remains sustained as long as the robot is not habituated to the object. Once the learning begins, the slow boost component periodically reset and delivers excitation to the tonic and phasic NF. The rate at which the boost reset depends on the persistence mechanism. In accordance with previous research [55] who suggests that the error prediction contributes to the phasic activation within the LC, our model propose an indirect contribution of the error through the persistence mechanism. Indeed, the persistence architecture modulates these periods depending on the goal error (see Appendix -D, Figures 19 and 20 - bottom row). If the error of a goal is significant, the period between two selections is lasting longer. At the first goal selection for the baseline experiment, the boost delivers excitation for 16 seconds before dropping down. At the end of learning the same skill, the boost provides excitation for 6.5 seconds. The transient action MT node also influences the period of boost activation. The $\tau_{+}$ parameter and the frequency of selecting a goal are inversely related. If $\tau_{+}$ is large, then the goal selection happens less often. In our model, the phasic activation within LC is the consequence of the goal selection for learning.
V-E A loop for incremental learning
Figure 16: Time course of object discovery and learning by introducing the cube first, the cylinder in second, and the ball last. Forward model error and learning progress depict the learning.
The architecture detects bottom-up events such as the presence of new objects in the scene and begins to explore through goal discovery. Here, we reduce the habituation time ($\tau_{+}$ in visual memory trace) so each object has a short exploration. We begin by introducing the cube, let the robot discover and learn a goal, then reproduce this step for the cylinder and the ball. This protocol exhibits the properties of the attentional system to focus on new stimuli. The time course of the experiment with the learning progress and forward model error is detailed in Figure 16. It is also possible to drive the robot into exploration again after habituation toward a specific location by increasing the object stimuli. In practice, it is sufficient to introduce a supplementary excitation within the object selection NF (habituation mechanism).
VI Conclusion and Discussion
The autonomous discovery and learning of goals in robotics is a complex issue to address and the state of the art represents a wide spectrum of approaches. Here, we take inspiration from neuroscience by rooting curiosity with attention in order to allow a robotic arm to discover and learn different skills. More precisely, we propose a simple model of the Locus Coeruleus, a nucleus in the brain controlling the switch between exploration and task engagement via the influence of curiosity, habituation and persistence. At first, the tonic activation within the LC drives the robot to generate actions that will produce new stimuli, and thus discover new goals via bottom-up attention. Then, the curiosity mechanism directly influences the Locus Coeruleus activation with a phasic mode by engaging in learning, depending on both forward model error and learning progress of these goals. In addition, we introduce new cognitive mechanisms to regulate the discovery and learning of goals with respectively a habituation and persistence component. The generation of robot motion is performed by inhibition of return and dynamic movement primitives. The complete architecture is modelled with dynamic neural fields (DNFs) except for the forward and inverse models, which are multi-layer perceptrons. The use of DNFs supports several technical novelties such as the generation of poses around an object and the interaction between skills in memory. In addition, the goal error modelling allows widespread connections that bootstrap the shift between goal discovery and learning as well as switch the learning of goals. Regarding the habituation paradigm driving exploration, we demonstrated that the robot was able to discover various goals, with a higher average number of goals for a slow habituation. For persistence, we tuned the time spent on learning a goal as well as the strength of the error inhibition to the LC and demonstrated a clear influence on learning. Varying these parameters manifests positive and negative impacts on learning that suggest an individual optimization depending on the objects. Moreover, the curiosity module exhibits an interesting property in case of two similar goals where the learning of one directly inhibits the other. Finally, the results indicate that the architecture continuously oscillates between goal discovery through bottom-up attention and goal learning via curiosity.
Nevertheless, the current model suffers from several limitations. The objects in habituation are distinguished base on colors, which is not realistic. In the future, we plan to use a variational autoencoder (VAE) to integrate more features and improve exploration. For the former, an object and a goal will be characterized by more features (3D object pose, touch sensor, rotation), so a more diverse set of skills can be learned. Regarding the latter, we will use the latent space as a supplementary method for discovering new goals through direct exploration [37]. This way, we will investigate the balance between direct exploration (VAE) and random exploration with the inhibition of return method. The end-effector orientation as well as the $z$ dimension of the robot were fixed and we intend to remove this limitation to perform a diverse range of robot actions that could lead to different outcomes and produce a developmental trajectory. Then, we will explore the tuning of the persistence and error inhibition parameters together. Indeed, there are several possibilities offered by the model and different learning patterns could still be observed. We would like to investigate if the number of goals discovered during habituation can be used to predict the parameters for an optimal learning. Finally, The goals discovered were on the same developmental level and the first learning selection entirely depended on the weights initialization of the forward models. We intend to investigate if the persistence mechanism can autonomously shift from a short to a long persistence in order to bootstrap the learning progress of the discovered goals.